Science Report – April 13th

Crew Scientist:

· Hydroponic garden completely assembled.

o Aeration system and pipes assembled and transported from RAM to the GreenHab

o Materials used from the RAM: gorilla glue, dewalt drill, hammer, razor blade, and needle-nosed pliers.

o Thermostat in the GreenHab set to 65 degrees F near 8:30pm

· Experimental steps completed today

o Calculated the amt of additives to be input into system.

o Administered additives (experimental variables: worm castings, BioLive fertilizers, compost biproducts)

· Troubleshooting

o Small leak at the ends of the pipes. Need to seal with caulking/ PVC pipe glue on Sol 4. Rate of evaporation may no longer be a feasible component to track unless leak can be sealed.

Crew Botanist:
Started to set up a botany experiment investigating the effect of different soil compositions on the growth of plants. The crew botanist recognized that there was a miscalculation for determining the group and population sizes. Namely, the weight fraction was used to calculate the quantities of different soil compositions instead of the volume fraction. If the problem was not corrected, the experimental group with increasing amount of martian regolith will decrease the overall volume of the "soil/regolith"; this would have meant that there would have been less material surrounding the seeds (and less potential nutrients). The crew botanist started to implement the fix but the completion of the work is rescheduled for tomorrow.

SH, XO
Red Planet People – MDRS Crew 245 "Team Patamars"To Mars and Beyond – For All!

EVA Report – April 13th

Crew 245 EVA Report 13-04-2021

EVA #2

Author: Shravan Hariharan, Crew Engineer

Purpose of EVA: Practice EVA to Pooh’s Corner, Round 2

Start time: 1100

End time: 1200

Narrative: Immediately after the first EVA Crew returned to the Habitat, it was time for the remaining three crew members to embark on their practice EVA! After discussing lessons learned with the initial EVA crew, the Commander, Crew Engineer, and Crew Botanist all suited up and embarked on EVA. The main goal of this Practice EVA was to formalize EVA protocols, including a medical check by the HSO prior to departure, and an order in which Crew Members should get ready for EVA. Upon arrival at Pooh’s Corner, the Crew practiced their Emergency Shelter Identification procedure – the first shelter was named the Dusty Ditch! After identification of the first shelter, both the Commander and Crew Engineer’s communications systems failed – the batteries of the radios were too low. At that point, it was determined that continuing on EVA was too dangerous, so the Crew Botanist communicated with the Hab, and led the crew back to the Habitat using hand signals. The Crew Commander is currently working on standardizing hand signals in case we have future communications failures. We have learned how to test the batteries prior to departing on EVA, so we are thankful that we were able to experience this failure and learn from it!

Destination: Pooh’s Corner

Coordinates (use UTM NAD27 CONUS): 12 S 0518875, UTM 4251175

Participants: Commander, Crew Engineer, Crew Botanist

Road(s) and routes per MDRS Map: Habitat to Entrance Road, left on Cow Dung Road, destination on right in ~0.5 km

Mode of travel: 1 Rover, 1 ATV


Shravan Hariharan, Crew Engineer
Red Planet People – MDRS Crew 245 "Team Patamars"
To Mars and Beyond – For All!

EVA Report – April 13th

Crew 245 EVA Report 13-04-2021

EVA #1

Author: Shravan Hariharan, Crew Engineer

Purpose of EVA: Practice EVA to Pooh’s Corner,

Start time: 0900

End time: 1030

Narrative: The three crew members were all extremely excited to participate in their first EVA, and take our first steps on Mars! First words were “We go to Mars together!” from our Executive Officer. The ATV started with a minor amount of choke and the rover worked flawlessly; the ride to Pooh’s Corner was relatively quick. XO observed that riding in the ATV with the two-piece suit could be uncomfortable. The MDRS map was accidentally left in the Habitat, so the remaining Crew Members in the Hab provided verbal navigation. Crew Scientist brought the crew to a location for soil sampling, and the three EVA members took photos and videos of each other to document the occasion. Three samples were collected while practicing the techniques and training one another, although the soil was rockier than expected. After taking some additional photos, the EVA crew drove back to the Habitat. Brief issues such as the Rover being left in “Forward” were quickly resolved, and the Crew returned to the Habitat. We are grateful to the Director for suggesting this practice EVA, as we were able to successfully practice our EVA protocols and operations, such as donning and doffing the suits, collecting supplies, and communicating as a team.

Destination: Pooh’s Corner

Coordinates (use UTM NAD27 CONUS): 12 S 0518875, UTM 4251175

Participants: Executive Officer, Crew Scientist, Health and Safety Officer

Road(s) and routes per MDRS Map: Habitat to Entrance Road, left on Cow Dung Road, destination on right in ~0.5 km

Mode of travel: 1 Rover, 1 ATV

Operations Report – April 13th

Crew 245 Operations Report 13-04-2021

SOL: 3

Name of person filing report: Shravan Hariharan

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator: Not run, as solar power was sufficient

Hours run: N/A

From what time last night: N/A

To what time this morning: N/A

List any additional daytime hours when the generator was run: N/A

Solar— SOC 80% (Before generator is run at night)

Diesel Reading – 50%

Station Propane Reading – ?% – Director, please confirm

Water (loft tank): 25 gallons

Water Meter: 150273.7 units

Water (static tank): ? gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Opportunity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: no, in town

Curiosity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Perseverance rover used: Yes

Hours: 200.3

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: 350.1

Reason for use: EVA

Oil Added? No

ATV Fuel Used: N/A

# Hours the ATVs were used today: 2.5 (but 0.5 hours of active transit)

Notes on ATVs: Mission Director will monitor and refuel ATVs for duration of mission

HabCar used and why, where?: No

CrewCar used and why, where?: No

General notes and comments: Hab kitchen sink is leaking, and seal will be re-caulked around border of sink. Leak is minor, and should not require any outside assistance.

Summary of internet: Crew has almost entirely consumed data for the day, so additional monitoring and rationing needed.

Summary of suits and radios: All radios nominal, suits currently charging. Suit batteries successfully swapped by Crew Engineer last night, and all suits functioned properly on EVAs today.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: None

Summary of ScienceDome operations: Used by Crew Botanist and Crew Scientist for botany experiments

Summary of RAM operations: Used by Crew Scientist to build hydroponics system

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: HSO report coming in with minor notes

Questions, concerns and requests to Mission Support: How do we check the amount of water in the static tank? We were unable to find a meter.


Shravan Hariharan, Crew Engineer
Red Planet People – MDRS Crew 245 "Team Patamars"
To Mars and Beyond – For All!

Crew Photos – April 13th

Having trouble with data for photos. Got these uploaded over our lunch hour, however do not expect to be able to get as many photos online. Discussed in the sol summary.

Shayna Hume, Executive Officer
Red Planet People – MDRS Crew 245 “Team Patamars”
To Mars and Beyond – For All!

Commander Report – April 13th

Entering Mars’ Atmosphere

Today, we woke up on Mars. We decided as a crew to enter sim overnight, and it was with an eerie knowledge that today would be very different from yesterday that I let my eyelids close. When I woke up – to the effervescent and ever-present chorus of Kanye’s “Good Morning” – it was with the jolting realization that as of today, I am joining the ranks of analog Martian astronauts.

We did not waste any time acclimating. This analog mission is only two weeks long, including every technology delay and equipment failure, which means that to accomplish the science we set out to do, we would need to fast-forward and immediately begin setting up shop.

A successful mission means that in simulation, science has been safely done. In order to do science, Martian astronauts need to do EVA’s, or extravehicular activities. It’s vital that EVA’s only be done when needed, as the science return has to justify the risk in venturing out onto the surface of another planet, but – as you can guess – we were all thrilled to have a reason to don our spacesuits.

The EVA

Five minutes. We stood in the airlock for five minutes, waiting for the go-ahead from the Hab. Once they gave us the say-so, it meant that the room we were huddled in had depressurized, and that we could open the sealed outer door and leave the Hab for the first time since the simulation began.

The first EVA comprised of our Executive Officer, Hume, HSO Officer, Coultrup, and Crew Scientist, Ettlin. We had only changed up our intended 3-person teams in order to make sure that the appropriate trainings were given during our practice EVA’s today. But – none of us minded the fact that it was our three female crewmembers who were the first to take that step. Our first words out on the surface were: “We go to Mars – together.” We care about diversity in the future of space, and it felt appropriate.

The EVA went wonderfully. There were a few small issues – bumping helmets together, having to squeeze into the rover with a bulky helmet and spacesuit on, and media taken being washed out by the harsh sunlight, but we returned with high spirits.

The second EVA – comprised of Commander Dickstein, Engineer Hariharan, and Botanist Hernandez started off enthusiastic and well. They received more challenges than the former team. Soon after leaving, Dickstein and Hariharan lost communications and Hernandez had to call back to the Hab. However, he quickly organized a plan for returning to the Hab using visual communications and hand signals. The team returned safely and calmly a few minutes later, handling the problem with ease and grace.

We all noted that because of the reality of the analog level, it truly felt as though we were on an expedition where our life support systems mattered and where we were preparing to do real science. We intend to begin our science EVA’s tomorrow, and in taking these trips today, were able to come up with not only EVA checklists, but also our best practices for traveling outside of the hab.

EVA Best Practices

· Charge your comms!

· Decide a nominal system of hand signals for events of communications outages

· Confirm all vehicles’ movement prior to leaving the Hab

· In the event of communications outages, do not lose visual on non-communicative members. The member with comms ought to follow up the rear in order to be able to call the Hab for safety reasons.

· Medium level is preferred for the air coolant system, as higher can result in comms being distorted.

· Maintaining space is important, as it’s easier to bump helmets than desired.

· Put a container of all needed supplies for EVA into the airlock prior to donning spacesuits – it’s easy to forget items after helmets are on.

· Customize a “preflight checklist” based on your crew’s needs.

· Name a “Hab controller” who will take point on all radio communications.

· Call out “Hab” when communications are asked from the Hab instead of the EVA members.

· Have an order of business and priorities for an EVA ahead of time in case of early return.

· More to come!

Shayna Hume, XO
Red Planet People – MDRS Crew 245 “Team Patamars”
To Mars and Beyond – For All!

Sol Summary – April 13th

Crew 245 Sol Summary Report 13-04-2021

Sol: 3

Summary Title: First day in Sim

Author’s name: Shayna Hume

Mission Status: Executive Officer

Sol Activity Summary:

· Woke up to “Good Morning” by Kanye…. again

· Early morning hours: Strawberry, blueberry, and chocolate chip pancakes! Chef Julio is unreal.

· Late morning hours: Hurried for a 9am first EVA. From 9-12, we did our two practice EVA’s in two crews. Crew 1: Hume, Coultrup, Ettlin. Crew 2: Dickstein, Hariharan, Hernandez. First words outside: “We go to Mars together.” Second EVA ended early due to comms breakdown.

· Noon: separated for a variety of activities. Crew Scientist doing botany. HSO starting a botany kit as well. Report-writing for the rest.

· Lunch: Chef Julio preparing quesadillas! Corn tortillas from scratch! It was absolutely wonderful.

· Afternoon: Scientist & Botanist in Greenhab, HSO and Commander working on Documentary, Engineer and XO on misc.

· Late Afternoon: Scientist & Botanist in Greenhab, Engineer grabbing people for dexterity task, HSO & XO answering misc. calls for help, Commander videoing intro videos for our mockumentary.

· Dinner: Beef tacos!

Look Ahead Plan: Our first science mission is tomorrow morning. In addition, we will be completing the second day of dexterity testing in spacesuit gloves. Misc. botany will be happening, as well as the first soil sample analysis for geology and astrobiology. We are now living fully in sim and beginning to work towards making our workload each day sustainable.

Anomalies in work: Comms went dead on EVA, ended early. Dead charge only.

Weather: Warm in daytime, around 40 F at nighttime. Clear weather. Expected rain did not fall.

Crew Physical Status: No major issues. HSO Report coming in for small aid kit use.

EVA: Two practice EVA’s. First one had several equipment issues, completed as nominal however. Second had comms issue. EVA Reports incoming.

Reports to be filed: HSO Report, Operations Report, EVA Report, EVA Plan, Science Report

Support Requested: We would like to abstain from sending 8 daily photos, as it’s taking up our whole internet. We uploaded photos over lunch and it virtually drained the entire internet for the day. A single photo may be more feasible. Please confirm a single photo of the day will do for now. We are running out just by sending this.

Shayna Hume, Executive Officer
Red Planet People – MDRS Crew 245 “Team Patamars”
To Mars and Beyond – For All!